OpenStudy (truwhovian11):

What is the distance between the points (15, 3) and (15, -9) in the xy-plane? 24 units 6 units 7.5 units 12 units

OpenStudy (anonymous):

Lets draw a number line here

OpenStudy (anonymous):

waaaaaaaaaaait

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (truwhovian11):

what this is middle school

hartnn (hartnn):

lets make it more simple, when the x co-ordinates are same, the distance is the difference between y co-ordinates. d = 3-(-9) = ... ?

OpenStudy (sparrow2):